Hypcon connectors


cmjones01
 

My 7104 is now working well, thanks to advice from this group. Even
the plugins are mostly OK: the 7B15 wouldn't trigger because no +5V
was getting to the trigger board due to a corroded connector. I've
given one 7A29 the epoxy treatment and it works well though needs the
low-frequency gain setting up. The other 7A29 has more serious
attenuator trouble and will have to wait.

Almost all the faults I've experienced have been due to the Hypcon
connectors not making proper contact. I guess that after 40 years
those elastomers are perhaps not as springy as they once were. What
I've found is that removing them and cleaning everything up and
carefully refitting doesn't usually help - I've done it three or four
times in some cases.

It seems that the elastomer part seems to acquire a shape which stops
some pins making contact. My strategy for getting the scope working
well has been to swap around the elastomers between different devices,
finding a combination for each one where all the pins that matter make
contact. To get both horizontal slots working properly I had to borrow
an elastomer from the poorly 7A29 to fit the horizontal channel
switch. The device I borrowed it from only uses four pins, so I'm
hoping that the less-than-perfect elastomer will still allow the 7A29
to work properly.

Apart from misbehaving Hypcons (and the expected scratchy pots and
switches) the only thing I've actually had to adjust was a bit of the
Z-axis calibration in order to get the range of the intensity
controls, especially the readout, right. There are indications from
labels I've found inside that this mainframe may have been sitting
unused since 1994, so it's done pretty well!

Chris